# Bellman Optimality of Average-Reward Robust Markov Decision Processes with a Constant Gain

### v1

@@ B.2 Proof of Theorem 4
Of course $E_{y}^{\Delta,q}[\tau_{y}]=0\leq|S|/\delta^{\prime}$. This implies Lemma 3.
Step 4: Equipped Lemma 3, we now proceed to analyze the span of $v_{\gamma}^{*}$. We consider bounding $v_{\gamma}^{*}(x)-v_{\gamma}^{*}(y)$ for a fixed pair of states $x\neq y\in S$.
we have that for all $s\in S$,
i.e. $\Delta_{\epsilon}$ is $\epsilon$-optimal. Moreover, there exists stationary $p_{\epsilon}\in\mathrm{K}_{\mathrm{S}}$ s.t.
Note that by the definition of $\Delta_{\epsilon}$ and $p_{\epsilon}$,
$v_{\gamma}^{*}(x)-v_{\gamma}^{*}(y)$ $=\sup_{\pi\in\Pi_{\mathrm{S}}}\inf_{\kappa\in\mathrm{K}_{\mathrm{H}}}v_{\gamma}^{\pi,\kappa}(x)-\sup_{\pi\in\Pi_{\mathrm{S}}}\inf_{\kappa\in\mathrm{K}_{\mathrm{H}}}v_{\gamma}^{\pi,\kappa}(y)$ (B.5)
$\leq\inf_{\kappa\in\mathrm{K}_{\mathrm{H}}}v_{\gamma}^{\Delta_{\epsilon},\kappa}(x)-\inf_{\kappa\in\mathrm{K}_{\mathrm{H}}}v_{\gamma}^{\Delta_{\epsilon},\kappa}(y)+\epsilon$
$\leq v_{\gamma}^{\Delta_{\epsilon},\kappa}(x)-v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y)+2\epsilon.$
for any $\kappa\in\mathrm{K}_{\mathrm{H}}$. Since $\epsilon>0$ can be arbitrarily small, it suffices to choose a $\kappa\in\mathrm{K}_{\mathrm{H}}$ (potentially depending on $\Delta_{\epsilon},p_{\epsilon},x,y$) so that $v_{\gamma}^{\Delta_{\epsilon},\kappa}(x)-v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y)$ is uniformly bounded.
This can be achieved using a similar argument as in Bartlett and Tewari, [1]. We consider a history-dependent adversary $\kappa=(\kappa_{0},\kappa_{1},\dots)\in\mathrm{K}_{\mathrm{H}}$ defined as follows. Let $g_{t-1}=(s_{0},a_{0},\dots,s_{t-1},a_{t-1})$ and
$\kappa_{t}(s^{\prime}|g_{t-1},s,a)=\begin{cases}q(s^{\prime}|s,a)&\text{if }s_{k}\neq y,\forall k\leq t-1\text{ and }s\neq y,\\ p_{\epsilon}(s^{\prime}|s,a)&\text{otherwise.}\end{cases}$ (B.6)
Here $q=q^{y,\Delta_{\epsilon}}$ is defined in Lemma 2 and 3. In other words, the $\kappa$ uses $q$ when the chain hasn’t hit $y$ and uses the $\epsilon$-optimal adversary after hitting $y$.
Under this history-dependent adversarial policy, we have
$v_{\gamma}^{\Delta_{\epsilon},\kappa}(x)$ $=E^{\Delta_{\epsilon},\kappa}_{x}\sum_{k=0}^{\infty}\gamma^{k}r(X_{k},A_{k})$ (B.7)
Note that by the construction of $\kappa$ in (B.6)
$E_{x}^{\Delta_{\epsilon},\kappa}\tau_{y}$ $=\sum_{k=0}^{\infty}P_{x}^{\Delta_{\epsilon},\kappa}(\tau_{y}\geq k)$
$=E_{x}^{\Delta_{\epsilon},q}\tau_{y}.$
Here, $(i)$ is becasue $x\neq y$, $(ii)$ follows from the definition of $E_{x}^{\Delta_{\epsilon},\kappa}$, $(iii)$ is because by (B.6)
and the last equality follows from reversing the previous steps.
On the other hand,
$E^{\Delta_{\epsilon},\kappa}_{x}\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})$ (B.8)
$=v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y).$
Here $(i)$ applies Fubini’s theorem leveraging the positivity of the summand, $(ii)$ follows from the definition of $\kappa$ in (B.6), $(iii)$ is obtained from reversing the previous equalities, and $(iv)$ is because $p_{\epsilon}$ is a stationary policy and, under $E_{x}^{\Delta_{\epsilon},p_{\epsilon}}$, $(X_{t},A_{t})$ is strong Markov.
Therefore, the bound in (B.5) and (B.7) implies that
$v_{\gamma}^{*}(x)-v_{\gamma}^{*}(y)$ $\leq v_{\gamma}^{\Delta_{\epsilon},\kappa}(x)-v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y)+2\epsilon$
$\leq E_{x}^{\Delta_{\epsilon},q}\tau_{y}+v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y)-v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y)+2\epsilon$
$\leq\frac{|S|}{\delta^{\prime}}+2\epsilon.$
where the last inequality follows from Lemma 3.
Since $\epsilon>0$, $x,y\in S$ are arbitrary, we have that
$\left|v_{\gamma}^{*}\right|_{\mathrm{span}}=\max_{x\in S}v_{\gamma}^{*}(x)-\min_{x\in S}v_{\gamma}^{*}(x)\leq\frac{|S|}{\delta^{\prime}}.$
So, $\left|v_{\gamma}^{*}\right|_{\mathrm{span}}$ is uniformly bounded in $\gamma$. By Theorem 3, we conclude that (3.1) has a solution. This completes the proof of Theorem 4. ∎
B.3 Proof of Theorem 5
@@
As in the proof of Theorem 4, we show that $\left|v_{\gamma}^{*}\right|_{\mathrm{span}}$ is uniformly bounded in $\gamma$. To achieve this, we will be considering the inf-sup version of the $\gamma$-discounted Bellman equation defined as follows:
Definition 5.
We say that $v^{\prime}_{\gamma}:S\rightarrow\mathbb{R}$ is the unique solution to the $\gamma$-discounted inf-sup equation if
$v_{\gamma}^{\prime}(s)=\inf_{p_{s}\in\mathcal{P}_{s}}\sup_{\phi\in\mathcal{Q}}E_{\phi,p_{s}}[r(s,A_{0})+\gamma v_{\gamma}^{\prime}(X_{1})],\quad\forall s\in S.$ (B.9)
Similar to (3.3), the existence and uniqueness of the solutions of (B.9) follow from a standard contraction argument as in Wang et al., 2023a [7].
We note that under the convexity assumptions in Theorem 5, if $v_{\gamma}^{\prime}$ is the solution to (B.9), then an application of Sion’s min-max theorem (see Corollary 2.1 in Wang et al., 2023a [7]) yields $v^{\prime}_{\gamma}=v^{*}_{\gamma}$. Therefore, it suffices to show that $\left|v_{\gamma}^{\prime}\right|_{\mathrm{span}}$ is uniformly bounded in $\gamma$.
The remainder of the proof follows closely the argument of Theorem 4. In particular, modified versions of the key lemmas (with the roles of the controller and adversary reversed) continue to hold. Therefore, we provide only an abbreviated argument here, emphasizing the differences.
Step 1: The following lemma, analogous to Lemma 1, holds:
@@
Under the assumptions of Theorem 5, there exists $\delta>0$ s.t. for any stationary adversary policy $p\in\mathcal{P}$ and any pair of states $x,y\in S$, there exists $\Delta$ and $N\leq|S|$ s.t. $p^{N}_{\Delta}(y|x)\geq\delta$.
We apply the same proof idea as before. For fixed $p\in\mathcal{P}$, by assumption, there exists $\Delta=\Delta^{x,y,p}\in\left\{{S\rightarrow\mathcal{Q}}\right\}$ so that $p_{\Delta}^{N}(y|x)>0$ for some $N=N^{x,y,p}$. Observe that for $\Delta,N$ fixed, the mapping $p\rightarrow p_{\Delta}^{N}(y|x)$ is continuous. So, there exists an open neighborhood $G_{p}$ of $p$ s.t. $q_{\Delta}^{N}(y|x)\geq p_{\Delta}^{N}(y|x)/2$ for all $q\in G_{p}$.
Since $\mathcal{P}$ is compact, by the same argument as in the proof of Lemma 1, there exists a finite set $C:=\left\{{p_{1},\dots,p_{|C|}}\right\}\subset\mathcal{P}$ s.t. $\left\{{G_{p}:p\in C}\right\}$ covers $\mathcal{P}$.
So, for any $q\in\mathcal{P}$, we can find $p\in C$ s.t. $q\in G_{p}$. Then, for any $x,y\in S$, set $\Delta=\Delta^{x,y,p}$ and $N=N^{x,y,p}$, we have
$q_{\Delta}^{N}(y|x)\geq\frac{1}{2}p_{\Delta}^{N}(y|x)\geq\min_{p\in C}p_{\Delta^{x,y,p}}^{N^{x,y,p}}(y|x)=:\delta(y|x)>0$
where $\delta(y|x)$ is independent of $p$. Let $\delta=\min_{x,y\in S}\delta(y|x)$, this implies Lemma 4.
Step 2: We show the following lemma.
@@
Under the assumptions of Theorem 5, there exists $\delta^{\prime}>0$ s.t. for any stationary adversary $p\in\mathcal{P}$ and states $y\in S$, there exists $\Delta:S\rightarrow\mathcal{Q}$ s.t.
$\min_{x\in S}P_{x}^{\Delta,p}(\tau_{y}\leq|S|)\geq\delta^{\prime}.$
For fixed adversary stationary policy $p\in\mathcal{P}$, we construct $\Delta=\Delta^{y,p}$ independent of $x$, that satisfies Lemma 5.
We will iteratively define $\Delta(\cdot|s)\in\mathcal{Q}$ until all $\left\{{\Delta(\cdot|s):s\in S}\right\}$ has been assigned. Denote $p_{\Delta}(\cdot|s)=\sum_{a\in A}\Delta(a|s_{i})p(\cdot|s,a)$.
We initialize the algorithm by assigning $\Delta(\cdot|y)=\phi(\cdot)$ for an arbitrary $\phi\in\mathcal{Q}$. Then, Let $V=\left\{{y}\right\}$ be the assigned states, and $V^{c}$ the compliment in $S$ are the unassigned states.
@@
    Choose any unassigned state $s_{0}\in V^{c}$. Then by Lemma 1 there exists $\xi=\xi^{x,y,p}:S\rightarrow\mathcal{Q}$ and $N$ s.t. $p_{\xi}^{N}(y|s_{0})\geq\delta$. Therefore, by the “trimming” argument in the proof of Lemma 2, there exists a non-repeating path $s_{0}\rightarrow s_{1}\rightarrow,\dots,\rightarrow s_{k}=y$, $s_{i}\neq s_{j}$ if $i\neq j$, s.t. $p_{\xi}(s_{i+1}|s_{i})\geq\delta|S|^{-N}$ for all $i\leq k-1$.
@@
    Let $j=\min\left\{{k\geq 1:s_{k}\in V}\right\}$ be the first index so that $s_{j}$ is assigned. So, $s_{i}\in V^{c}$ for all $i\leq j-1$. We assign $\Delta(\cdot|s_{i}):=\xi^{s_{0},y,p}(\cdot|s_{i})\in\mathcal{Q}$.
    Note that, at the current iteration, $\Delta(\cdot|s)$ is well-defined for all $s\in V$. Since $s_{j}\in V$, there is a non-repeating path $\left\{{s_{j}=s_{j}^{\prime},s^{\prime}_{j+1},\dots,s^{\prime}_{k}=y}\right\}\subset V$ s.t. $p_{\Delta}(s^{\prime}_{i+1}|s^{\prime}_{i})\geq\delta|S|^{-N}$.
    Therefore, after assigning $\Delta(\cdot|s_{i})$ for $i\leq j-1$, we have a new path $s_{0}\rightarrow\dots\rightarrow s_{j}\rightarrow s_{j+1}^{\prime}\rightarrow\dots\rightarrow s_{k}^{\prime}=y$ with positive one step transition probabilities at least $\delta|S|^{-N}$ under $p_{\Delta}$. We record this path that leads to $y$.
@@
    Update $V\leftarrow V\cup\left\{{s_{0},\dots,s_{j-1}}\right\}$.
Iterate until $V=S$.
@@
This algorithm terminates in at most $|S|$ iterations, producing $\Delta:S\rightarrow\mathcal{Q}$. Moreover, it produces a directed graph whose edges correspond to positive transition probabilities at least $\delta|S|^{-N}$ under $p_{\Delta}(\cdot|\cdot)$, ensuring that every state can reach $y$ in at most $|S|$ steps.
Therefore, we conclude that for any $y\in S$ and $p\in\mathcal{P}$, let $\Delta$ be constructed by the above algorithm, then
$\min_{x\in S}P_{x}^{\Delta,p}(\tau_{y}\leq|S|)\geq\left(\delta|S|^{-N}\right)^{|S|}=:\delta^{\prime}>0.$
Note that $\delta^{\prime}$ is independent of $x,y,\Delta$. This shows Lemma 5.
Step 3: The same proof as that of Lemma 3 implies the following Lemma.
Lemma 6.
Under the assumptions of Theorem 5, for any $p\in\mathcal{P}$ and $y\in S$, the $\Delta$ constructed in Lemma 5 satisfies
$\max_{x\in S}E^{\Delta,p}_{x}\tau_{y}\leq\frac{|S|}{\delta^{\prime}}.$
Step 4: By convexity of $\mathcal{Q}$ and $\mathcal{P}_{s}$ and the compactness of $\mathcal{P}_{s}$, Corollary 2.1 in Wang et al., 2023a [7] implies the existence of $\epsilon$-optimal stationary strategies $\Delta:S\rightarrow\mathcal{Q}$ and $p_{\epsilon}\in\mathcal{P}$ of both the controller and the adversary. In particular,
@@
$\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}v^{\Delta_{\epsilon},\kappa}_{\gamma}(s)+\epsilon\geq v^{*}_{\gamma}(s)=v^{\prime}_{\gamma}(s)\geq\sup_{\pi\in\Pi_{\mathrm{H}}}v^{\pi,p_{\epsilon}}_{\gamma}(s)-\epsilon.$
Moreover, observe that
$v^{\Delta_{\epsilon},\kappa_{\epsilon}}_{\gamma}(s)+\epsilon\geq\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}v^{\Delta_{\epsilon},\kappa}_{\gamma}(s)+\epsilon\geq v_{\gamma}^{*}(s)\geq\sup_{\pi\in\Pi_{\mathrm{H}}}v^{\pi,p_{\epsilon}}_{\gamma}(s)-\epsilon\geq v^{\Delta_{\epsilon},\kappa_{\epsilon}}_{\gamma}(s)-\epsilon.$
So,
$|v_{\gamma}^{*}(s)-v^{\Delta_{\epsilon},\kappa_{\epsilon}}_{\gamma}(s)|\leq\epsilon.$ (B.10)
Next, following the same proof, we construct a two-phase history dependent controller policy by first using $\Delta$ constructed in 5 and 6 until the first time hitting $y$ and then use an $\epsilon$-optimal stationary policy $\Delta_{\epsilon}$ afterwards. Let us denote this history-dependent controller by $\pi$.
By convexity of $\mathcal{Q}$ and $\mathcal{P}_{s}$ and the compactness of $\mathcal{P}_{s}$, Corollary 2.1 of Wang et al., 2023a [7] implies that
$v_{\gamma}^{\prime}(x)-v_{\gamma}^{\prime}(y)$ $=\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\sup_{\pi\in\Pi_{\mathrm{H}}}v_{\gamma}^{\pi,p}(x)-\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\sup_{\pi\in\Pi_{\mathrm{H}}}v_{\gamma}^{\pi,\kappa}(y)$ (B.11)
$\geq\sup_{\pi\in\Pi_{\mathrm{H}}}v_{\gamma}^{\pi,p_{\epsilon}}(x)-\sup_{\pi\in\Pi_{\mathrm{H}}}v_{\gamma}^{\pi,p_{\epsilon}}(y)-\epsilon$
$\geq v_{\gamma}^{\pi,p_{\epsilon}}(x)-v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(x)-3\epsilon,$
where the last inequality follows from (B.10).
We observe that
$v_{\gamma}^{\pi,p_{\epsilon}}(x)$ $=E^{\pi,p_{\epsilon}}_{x}\sum_{k=0}^{\infty}\gamma^{k}r(X_{k},A_{k})$
$=E^{\pi,p_{\epsilon}}_{x}\sum_{k=0}^{\tau_{y}-1}\gamma^{k}r(X_{k},A_{k})+E^{\pi,p_{\epsilon}}_{x}\gamma^{\tau_{y}}\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})$
$\geq E^{\pi,p_{\epsilon}}_{x}(\gamma^{\tau_{y}}-1)\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})+E^{\pi,p_{\epsilon}}_{x}\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k}).$
Note that
$E^{\pi,p_{\epsilon}}_{x}(\gamma^{\tau_{y}}-1)\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})$ $\geq E^{\pi,p_{\epsilon}}_{x}\frac{\gamma^{\tau_{y}}-1}{1-\gamma}$
$=-E^{\pi,p_{\epsilon}}_{x}\sum_{k=0}^{\tau_{y}-1}\gamma^{k}$
$\geq-E_{x}^{\pi,p_{\epsilon}}\tau_{y}$
$=-E_{x}^{\Delta,p_{\epsilon}}\tau_{y}$
$\geq-\frac{|S|}{\delta^{\prime}}$
and by (B.8),
$E^{\pi,p_{\epsilon}}_{x}\sum_{k=\tau_{y}}^{\infty}\gamma^{k-\tau_{y}}r(X_{k},A_{k})=v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(y).$
Therefore, from (B.11), we have that
$v_{\gamma}^{\prime}(x)-v_{\gamma}^{\prime}(y)\geq-\frac{|S|}{\delta^{\prime}}-3\epsilon.$
This hold for all $x,y\in S$, and $\epsilon$ can be arbitrarily small. Hence, we have that $|v_{\gamma}^{\prime}(x)-v_{\gamma}^{\prime}(y)|\leq|S|/\delta^{\prime}$ for all $x,y\in S$.
Under the assumptions of Theorem 5, Corollary 2.1 in Wang et al., 2023a [7] implies that $v_{\gamma}^{\prime}=v_{\gamma}^{*}$ solves (3.3) and (B.9). So, we conclude that

### v2

@@ C.3 Proof of Lemma 4
Of course $E_{y}^{\Delta,q}[\tau_{y}]=0\leq|S|/\delta^{\prime}$. This implies Lemma 4. ∎
Appendix D Proof of Theorem 6
The proof follows arguments very similar to those in Theorem 5. To avoid excessive repetition, we focus on explaining how the earlier proof carries over and highlighting the necessary modifications.
As in the proof of Theorem 5, we primarily address the weakly communicating case in assumption (1). Since a communicating adversary is a special case, the argument under assumption (2) carries over with only minor changes, the main difference being the lack of convexity of $\mathcal{Q}$ relative to assumption (1).
@@ Appendix D Proof of Theorem 6
From the argument proving Theorem 4, it follows that if the solutions $\left\{{v^{\prime}_{\gamma}:\gamma\in(0,1)}\right\}$ to (3.4) have uniformly bounded span, then (3.2) has a solution. Therefore, we now proceed to establish the uniform boundedness of $\left|v^{\prime}_{\gamma}\right|_{\mathrm{span}}$, mirroring the argument used in the proof of Theorem 5.
Preliminary Constructions
Mirroring the proof of Lemma 1, it is easy to see that the following Lemma holds.
@@
If the stationary adversary policy $p\in\mathcal{P}$ is weakly communicating and $\mathcal{P}$ is S-rectangular, then for any $w\in S$ and any $y\in C_{p}$, there exist $\Delta:S\to\mathcal{Q}$ and $N\leq|S|$ such that $p_{\Delta}^{N}(y|w)>0$.
Next, we will leverage the compactness of $\mathcal{P}$ to construct a finite subset $B^{\prime}$ of stationary adversary policies that yields a uniform lower bound on hitting probabilities.
Assume the adversary is weakly communicating, $\mathcal{P}$ is S-rectangular, and $\mathcal{Q}$ and $\mathcal{P}$ are compact. By Lemma 5, for any stationary policy $p\in\mathcal{P}$ and any $w\in S$, $y\in C_{p}$, there exist $\Delta$ and $N\leq|S|$ (both possibly depending on $(w,y,p)$) such that $p_{\Delta}^{N}(y|w)>0$.
Note that if we fix $\Delta=\Delta^{w,y,p}$ and $N=N^{w,y,p}$ given by Lemma 5, then the mapping $q\rightarrow q_{\Delta}^{N}(y|w)$ is continuous in $q$. So, there must exist an open neighborhood $O_{p}\subseteq\mathcal{P}$ of $p$ such that
$q_{\Delta}^{N}(y|w)\geq\tfrac{1}{2}p_{\Delta}^{N}(y|w),\quad\forall q\in\overline{O_{p}}.$ (D.1)
We also consider another (not necessarily open) cover $\left\{{K_{p}:p\in\mathcal{P}}\right\}$ of $\mathcal{P}$. For any fixed $p\in\mathcal{P}$, if $C_{p}=S$, i.e. the entire state space is communicating, then let $K_{p}=\mathcal{P}$. On the other hand, if $C_{p}^{c}\neq\varnothing$, we construct $K_{p}$ as in the following lemma.
@@
Assume the adversary is weakly communicating, $\mathcal{P}$ is S-rectangular, and $\mathcal{Q}$ and $\mathcal{P}$ are compact. Then, for each $p\in\mathcal{P}$ with $C_{p}^{c}\neq\varnothing$, there exists an open neighborhood $K_{p}$ of $p$ such that
$0\leq\sup_{\Delta:S\to\mathcal{Q}}e^{\top}(I-M^{p}_{\Delta,q})^{-1}e\leq\sup_{\Delta:S\to\mathcal{Q}}e^{\top}(I-M^{p}_{\Delta,p})^{-1}e+1<\infty,\quad\forall q\in\overline{K_{p}},$ (D.2)
where both suprema are attained. In this expression, $M^{p}_{\Delta,q}$ is the principal submatrix of $q_{\Delta}$ on $C^{c}_{p}$ defined by $M^{p}_{\Delta,q}(s,s^{\prime})=q_{\Delta}(s^{\prime}|s)$ for $s,s^{\prime}\in C^{c}_{p}$, and $e$ denotes the all-ones vector in $\mathbb{R}^{|C_{p}^{c}|}$.
Remarks on the proof of Lemma 6.
The proof of Lemma 6 follows from the same argument as that of Lemma 2. In particular, the invertibility of $I-M_{\Delta,p}^{p}$ follows from the definition of $M_{\Delta,p}^{p}$ being the transient part of $p_{\Delta}$; the continuity of $(\eta,q)\rightarrow(I-M_{\eta,q}^{p})^{-1}$ within $\left\{{S\rightarrow\mathcal{Q}}\right\}\times K^{\prime}_{p}$, where $K_{p}^{\prime}$ is an open neighborhood of $p$, follows from the continuity of
$q\rightarrow\min_{\Delta:S\rightarrow\mathcal{Q}}\left|\det(I-M_{\Delta,q}^{p})\right|$
and the invertibility of $I-M_{\Delta,p}^{p}$ for all $\Delta:S\rightarrow\mathcal{Q}$.
These properties imply the finiteness and continuity of
$q\rightarrow\max_{\Delta:S\rightarrow\mathcal{Q}}e^{\top}(I-M_{\Delta,q}^{p})^{-1}e$
within some open neighborhood of $p$, implying Lemma 6. ∎
As in the proof of Theorem 5, we have defined $O_{p}$ and $K_{p}$ for all $p\in\mathcal{P}$. With these constructions, we define
$G_{p}:=O_{p}\cap K_{p}.$
Note that when $C_{p}^{c}=\varnothing$, then $G_{p}=O_{p}\ni p$ is non-empty and open. When $C_{p}^{c}\neq\varnothing$, both $O_{p}$ and $K_{p}$ are open neighborhoods of $p$. Therefore, $\left\{{G_{p}:p\in\mathcal{P}}\right\}$ is an open cover of $\mathcal{P}$. Since $\mathcal{P}$ is compact, there exists a finite subcover $\left\{{G_{p}:p\in B^{\prime}}\right\}$ where $B^{\prime}:=\left\{{p_{1},\ldots,p_{|B^{\prime}|}}\right\}\subseteq\mathcal{P}$ is a finite subset.
With this construction, it is clear that the proof of Lemma 7 carries over, and we have the following result.
Lemma 7.
Under the assumptions of Theorem 6, there exists $\delta>0$ such that, for any stationary adversary policy $p\in\mathcal{P}$ with $p\in G_{p_{k}}$ and any $y\in C_{p_{k}},w\in S$, there exists $\Delta:S\to\mathcal{Q}$ and $N\leq|S|$ such that $p_{\Delta}^{N}(y|w)\geq\delta$.
Again, note that following the same argument in Remark 3, if we replace $G_{p_{k}}$ with $O_{p_{k}}$, we do not need the compactness of $\mathcal{Q}$ to show Lemma 7. Moreover, as Lemma 3 implies Lemma 4, Lemma 7 implies the following expected hitting time bound.
Lemma 8.
Under Lemma 7, there exists $\delta^{\prime}>0$ s.t. for any stationary adversary policy $p\in\mathcal{P}$ with $p\in G_{p_{k}}$ and $y\in C_{p_{k}}$, there exists $\Delta:S\to\mathcal{Q}$ such that
$\max_{w\in S}E^{\Delta,p}_{w}\tau_{y}\leq\frac{|S|}{\delta^{\prime}},$
where $\tau_{y}=\inf\left\{{t\geq 0:X_{t}=y}\right\}$.
Proof of Lemma 8 based on the proof of Lemma 4.
Fix $p\in\mathcal{P}$ and $y\in S$. By Lemma 7, for any $w\in S$ there exist a stationary controller $\Delta:S\rightarrow\mathcal{Q}$ and $N\leq|S|$ such that
$P^{\Delta,p}_{w}(\tau_{y}\leq|S|)\geq p_{\Delta}^{N}(y|w)\geq\delta.$
As in the proof of Lemma 4, it sufficies to show that for some $\delta^{\prime}>0$ independent of $y$, there is a controller $\eta:S\rightarrow\mathcal{Q}$ such that
$\min_{w\in S}P_{w}^{\eta,p}(\tau_{y}\leq|S|)\geq\delta^{\prime}.$
That is, the choice of controller can be made independent of the initial state $w$.
To this end, mirroring the proof of Lemma 4, let us define $\eta=\eta^{y,p}$ algorithmically as follows.
We initialize by assigning $\eta(\cdot|y)=\phi$ for an arbitrary $\phi\in\mathcal{Q}$. Then let $V=\left\{{y}\right\}$ be the assigned states, and $V^{c}=S\setminus V$ the unassigned states.
@@
    Choose any unassigned state $s_{0}\in V^{c}$. By Lemma 7, there exists $\Delta:S\rightarrow\mathcal{Q}$ and $N\leq|S|$ such that $p_{\Delta}^{N}(y|s_{0})\geq\delta$. Following the argument in the proof of Lemma 4, this implies that there exists a non-repeating path $s_{0}\rightarrow s_{1}\rightarrow\dots\rightarrow s_{k}=y$ with $k\leq N$ and $p_{\Delta}(s_{i+1}|s_{i})\geq\delta|S|^{-N}$ for $i=0,\dots,k-1$.
@@
    Let $j=\min\{i\geq 1:s_{i}\in V\}$ be the first index on the path already in $V$. For $i=0,\dots,j-1$ set $\eta(\cdot|s_{i}):=\Delta(\cdot|s_{i})\in\mathcal{Q}$. Then $p_{\eta}(s_{i+1}|s_{i})\geq\delta|S|^{-N}$ for all such $i$. Moreover, since $s_{j}\in V$, there is a path from $s_{j}$ to $y$ already recorded with each edge at least $\delta|S|^{-N}$. Concatenating, we obtain a path from $s_{0}$ to $y$ with all edge probabilities bounded below by $\delta|S|^{-N}$. Record this path.
@@
    Update $V\leftarrow V\cup\{s_{0},\dots,s_{j-1}\}$.
Repeat until $V=S$.
@@
The algorithm terminates in at most $|S|$ iterations, producing $\eta\in\{S\to\mathcal{Q}\}$, since we always assign $\eta(\cdot|s)\in\mathcal{Q}$. Moreover, it produces a directed graph whose edges correspond to positive transition probabilities at least $\delta|S|^{-N}$ under $p_{\eta}(\cdot|\cdot)$, ensuring that every state can reach $y$ in at most $|S|$ steps. Therefore,
$\min_{w\in S}P_{w}^{\eta,p}(\tau_{y}\leq|S|)\geq(\delta|S|^{-N})^{|S|}=:\delta^{\prime}>0.$
This and the Markov renewal argument in the proof of Lemma 4 imply Lemma 8. ∎
Decomposing $\left|v^{\prime}_{\gamma}\right|_{\mathrm{span}}$
First, note that $v^{\prime}_{\gamma}$ solves (3.4). For each $\epsilon>0$, there exist $p_{\epsilon}\in\mathcal{P}$ that is $\epsilon$-optimal in the following sense
@@
$v^{\prime}_{\gamma}(s)=\inf_{p\in\mathcal{P}}\sup_{\pi\in\Pi_{\mathrm{H}}}v_{\gamma}^{\pi,\kappa}(s)\leq\sup_{\pi\in\Pi_{\mathrm{H}}}v_{\gamma}^{\pi,p_{\epsilon}}(s)+\epsilon,\quad\forall s\in S.$
Since $\left\{{G_{p}:p\in B^{\prime}}\right\}$ covers $\mathcal{P}$, there is $p_{k}\in B^{\prime}$ s.t. $p_{\epsilon}\in G_{p_{k}}$.
Moreover, by the stationary optimality of MDPs, there exists $\Delta_{\epsilon}:S\rightarrow\mathcal{Q}$ such that
$\sup_{\pi\in\Pi_{\mathrm{H}}}v_{\gamma}^{\pi,p_{\epsilon}}(s)\leq v_{\gamma}^{\Delta_{\epsilon},p_{\epsilon}}(s)+\epsilon,\quad\forall s\in S.$
Similar to the decomposition for $\left|v^{*}_{\gamma}\right|_{\mathrm{span}}$, we let
$s_{\vee}\in\underset{s\in S}{\operatorname{arg}\,\operatorname{max}}\;v_{\gamma}^{\prime}(s),\quad s_{\wedge}\in\underset{s\in S}{\operatorname{arg}\,\operatorname{min}}\;v^{\prime}_{\gamma}(s),\quad\text{and}\quad y_{\vee}\in\underset{s\in G_{p_{k}}}{\operatorname{arg}\,\operatorname{max}}\;v^{\prime}_{\gamma}(s).$